OpenStudy (anonymous):

Check three of my answers to three different problems? Question #1: Evaluate x-3 for x = 2 My answer: 1/8. Question #2: Simplify 2a0. My Answer: 2. Question #3: Simplify (5/9)^- 2. My Answer: So here we can use the negative to flip around the fraction. So once we flip it around we get (9/5)^2. So the negative is gone, but the two is still there. So now we we can solve this with 9/5 * 9/5. We can multiply straight across for the top and get 81, for the bottom we multiply straight across and get 25. So the answer is 81/25.

OpenStudy (phi):

yes, all three are correct. if you mean x^-3 (it looks like x-3 or 2-3= -1. you should put in the ^ to show the -3 is an exponent) 2^-3 is 1/2^3 = 1/(2*2*2)= 1/8 same for 2a^0 where a^0 is 1 (anything to the 0 power is 1, except for 0^0 which is not defined) your explanations #3 is excellent

OpenStudy (anonymous):

Thanks.